HUBBLE SPACE TELESCOPE DAILY REPORT #2108 PERIOD COVERED: 0000Z (UTC) 04/22/98 - 0000Z (UTC) 04/23/98 Daily Status Report as of 113/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled with no anomalies. 1.3 Completed FOC/96/WF/PC-2/NIC/1/2/3/STIS/CCD 7591 (1997 Spring Heptathlon) The Faint Object Camera (f/96), the WF/PC-2, the Near Infrared Camera (Cameras One, Two, and Three), and the Space Telescope Infrared Spectrograph were used to make calibration observations as part of the Spring Heptathlon. This test is a continuation of the focal plane monitoring program. The goal is to measure relative positions of the SIs to within 50 milli-arcseconds, to verify the focal plane plate scale, and by repetition, to assess stability of the focal plane and FGSs. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-3627 and NGC-2903 far from the center by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The proposal completed as expected. 1.5 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.6 Completed NIC/1/2/3 7505 (The Hubble Diagram for Distant Supernovae Measuring Cosmic Deceleration and Global Curvature) The Near Infrared Cmaera (Cameras One, Two and Three) was used to make observations of Super Nova SN1998E in order to produce a precise Hubble diagram from measuring cosmic deceleration and global curvature. HST can make an essential contribution by providing photometry with substantially higher precision and better temporal coverage than can be obtained from the ground. The observations were completed as planned, and no anomalies were reported. 1.7 Completed Three Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.8 Completed NIC/1/2/3 7218 (Nearby Starburst Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the nearby starbursting galaxy M-82. This proposal encompasses a study of both the stars and interstellar material in starburst nuclei which will test ideas about the triggering of starbursts and whether starbursts propagate outwards in a nucleus. As described in 2.1, one of the re-acquisitions for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as planned, and no anomalies were reported. 1.9 Completed Three Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.10 Completed STIS/CCD 7635 (STIS CCD Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to test the hot pixel annealing procedure. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 The acquisition at 113/041048Z defaulted to fine lock backup on FGS-3 only when FGS-1 experienced a walkdown failure. HSTAR 6446 was written. Three proposals which will be dwscribed in tomorrow's report may have been affected: WF/PC-2 7909, NICMOS 7596, and STIS 7510. Scheduled Reacquisitions : 11 Successful : 11 The re-acquisition at 112/123216Z defaulted to fine lock backup on FGS-1 only when the FGS-3 scan step limit was exceeded. HSTAR 6445 was written. The proposal described in 1.8 may have been affected. 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Operations Notes: The NSSC-1 status buffer was dumped at 112/1401Z using ROP NS-3. Per ROP SR-1, the SSR correctable EDAC error counter was cleared on two occasions. The Battery 3 capacity check test continues. Discharge was terminated nominally at 15 Volts by the special hardware limit test at 112/1254Z. Engineering data recording for the test was stopped at 112/1311Z, and was subsequently played back. Recharge was started at 112/1432Z, the normal orbit day entry. The charge acceptance behavior is typical relative to that of other batteries. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH